OpenStudy (anonymous):

Given the sum of the series, Find the value of d. Please Help. Will give medal and fan :-)

OpenStudy (anonymous):

\[3+\frac{1}{4}(3+d)+\frac{1}{4^2}(3+2d)+....\infty = 8\]

OpenStudy (anonymous):

@ganeshie8 @ParthKohli @experimentX

OpenStudy (anonymous):

@e.mccormick @eliassaab @AccessDenied @beccaboo333 @Chamberlain15 @david111 @Evictu_FB @fratlife19

OpenStudy (anonymous):

@rational @mathmale

OpenStudy (experimentx):

collect 1/4s ... and d/4's

OpenStudy (anonymous):

@experimentX : I do understand that this particular thing deals with AGP ..

OpenStudy (experimentx):

just get rid of 'd' by taking common at the moment.

OpenStudy (anonymous):

But that's what we're supposed to find! 'd'!

OpenStudy (anonymous):

@ganeshie8 @Hero @hartnn @dumbcow @Nurali @nincompoop @eliassaab @AustinC @CO_oLBoY @ranga @tyr1 @UH60blackhawk @Valestrum @xTribalKing93x @zzr0ck3r

OpenStudy (shiraz14):

@newtoos: @dumbcow is correct. Please use the formula provided for S(inf) of the AGP as follows: \[S_{\infty} = \frac{ a }{ 1-r } + \frac{ rd }{ (1-r)^{2} }\], where S(inf)=8, a=3 & r=1/4.

OpenStudy (anonymous):

Shoot. I knew that formula. i know how to use it. I knew the a. But the whole time i took r as 4 not 1/4 :'(